# Solving Large Multicommodity Network Flow Problems on GPUs

### v1

@@ 1.1 Multicommodity network flow problem
Following [YDLB19] we aggregate all flows with the same destination, considering it to be one commodity that is conserved at all nodes except the source and destination, but can be split and combined. The commodity flows are given by the (destination-based) flow matrix $F\in{\mbox{\bf R}}^{n\times m}$, where $F_{i\ell}\geq 0$ denotes the flow on edge $\ell$ that is destined to node $i$. The edge capacity constraint can be expressed as $F^{T}\mathbf{1}\leq c$, where the inequality is elementwise.
@@
The MCF problem (1) always has solution. To see that it is always feasible, consider a unit flow from each source to each destination, over the shortest path, i.e., smallest number of edges, which exists since the graph is completely connected. We denote this flow matrix as $F^{\text{sp}}$. Now take $F=\alpha F^{\text{sp}}$, where $\alpha=1/\max_{i}(({F^{\text{sp}}}^{T}\mathbf{1})_{i}/c_{i})$, so we have $F^{T}\mathbf{1}\leq c$. Evidently $F$ is feasible, and we have $T_{ij}=\alpha$ for $i\neq j$. Existence of a solution follows since feasible set is bounded. The solution need not be unique. The optimal $T$, however, is unique.
@@ 1.2 Optimality condition and residual
A flow matrix $F$ is optimal for (2) if and only if $F\in\mathcal{F}$, $-FA^{T}\in\mathcal{T}$, and for all $Z\in\mathcal{F}$,
$\mathop{\bf Tr}(Z-F)^{T}G\geq 0,$
where $G=\nabla_{F}(-U)(-FA^{T})$ (see, e.g., [BV04, §4.2.3]). We have $G=U^{\prime}A$, where $U^{\prime}_{ij}=u_{ij}^{\prime}((-FA^{T})_{ij})$.
We can express this optimality condition in terms of projection onto the feasible flow set. Let $\Pi$ denote Euclidean projection onto $\mathcal{F}$. The optimality condition can be expressed in terms of a matrix $Q\in{\mbox{\bf R}}^{n\times m}$, with $F=\Pi(Q)$. Evidently $F\in\mathcal{F}$. We need to have $-FA^{T}\in\mathcal{T}$ for the gradient $G$ to exist. When that holds, the other optimality condition can be expressed as
$G=\gamma(F-Q)$
for some $\gamma\geq 0$.
@@ 1.3 Related work
Specialized to GPU-based optimizers for network flow optimization, [WZR+18] considers implementing a parallel routing algorithm on GPUs for SDN networks, which solves the Lagrangian relaxation of a mixed integer linear program. [KOY+15] implements a genetic method on GPUs for solving an integer linear program formulation of routing problem. [ZAC23] considers a linear program formulation of multicommodity network flow problems and constructs a deep learning model for generating new columns in delayed column generation method. [WHH12] implements an asynchronous push-relabel algorithm for single commodity maximum network flow problem, which is CPU-GPU hybrid. [LHQ+24] exploits exactly the same flow aggregation formulation of MCF following [YDLB19] as we do and trains a neural network model for minimizing unconstrained Lagrangian relaxation objective, and feeds the result as warm start to Gurobi [Gur24] to get the final answer.
@@ Acknowledgements
We thank Anthony Degleris and Parth Nobel for valuable discussions on implementation details.

### v2

@@ 1.1 Multicommodity network flow problem
Following [YDLB19] we aggregate all flows with the same destination, considering it to be one commodity that is conserved at all nodes except the source and destination, but can be split and combined. The commodity flows are given by the (destination-based) flow matrix $F\in{\mbox{\bf R}}^{n\times m}$, where $F_{i\ell}\geq 0$ denotes the flow on edge $\ell$ that is destined to node $i$. The edge capacity constraint can be expressed as $F^{T}\mathbf{1}\leq c$, where the inequality is elementwise. A similar flow aggregation formulation, though source-based, was considered in [BGB03].
@@
First let us show the MCF problem (1) is always feasible. Consider a unit flow from each source to each destination, over the shortest path, i.e., smallest number of edges, which exists since the graph is completely connected. We denote this flow matrix as $F^{\text{sp}}$. Now take $F=\alpha F^{\text{sp}}$, where $\alpha=1/\max_{\ell}(({F^{\text{sp}}}^{T}\mathbf{1})_{\ell}/c_{\ell})>0$, so we have $F^{T}\mathbf{1}\leq c$. Evidently $F$ is feasible, and we have $T_{ij}=\alpha>0$ for $i\neq j$, so $T=-FA^{T}\in\mathcal{T}$. This shows that the problem is always feasible. Let $U^{\text{sp}}$ denote the corresponding objective function.
We can add the constraint $U(T)\geq U^{\text{sp}}$ to the problem, without changing the solution set. With this addition, the feasible set is compact. It follows that the MCF problem (1) always has solution. The solution need not be unique. The optimal $T$, however, is unique. We also note that the argument above tells us that the implicit constraint $T=-FA^{T}\in\mathcal{T}$ is redundant.
@@ 1.2 Optimality condition and residual
Let $\tilde{\mathcal{F}}$ denote the closure of the feasible set, including the implicit constraint $T=-FA^{T}\in\mathcal{T}$,
~=$\tilde{\mathcal{F}}=\mathcal{F}\cap\{F\mid-FA^{T}\in\textbf{cl}(\mathcal{T})\},$
where $\textbf{cl}(\mathcal{T})$ denotes the closure of $\mathcal{T}$.
Then $F$ is optimal for (2) if and only if $F\in\mathcal{F}$, $-FA^{T}\in\mathcal{T}$, and
$\mathop{\bf Tr}(Z-F)^{T}G\geq 0$
holds for all $Z\in\tilde{\mathcal{F}}$, where $G=\nabla_{F}(-U)(-FA^{T})$ (see, e.g., [BV04, §4.2.3]). We have $G=U^{\prime}A$, where $U^{\prime}_{ij}=u_{ij}^{\prime}((-FA^{T})_{ij})$.
Optimality condition via projection onto $\mathcal{F}$.
For future use, we express the above optimality condition in terms of projection of a matrix $Q$ onto $\mathcal{F}$. Let $\Pi$ denote Euclidean projection onto $\mathcal{F}$. Suppose $Q\in{\mbox{\bf R}}^{n\times m}$, and set $F=\Pi(Q)$, so $F\in\mathcal{F}$. Suppose in addition that $-FA^{T}\in\mathcal{T}$, so that $G=\nabla_{F}((-U)(-FA^{T})$ exists. Then $F$ is also Euclidean projection of $Q$ onto $\tilde{\mathcal{F}}$. It follows that $\mathop{\bf Tr}(Z-F)^{T}G\geq 0$ for all $Z\in\tilde{\mathcal{F}}$, so the optimality condition above holds, and $F$ is optimal. Evidently it would hold if the weaker condition
$G=\gamma(F-Q)~{}\mbox{for some}~{}\gamma\geq 0$
holds.
Summarizing: $F$ is optimal if $F=\Pi(Q)$ for some $Q$, $-FA^{T}\in\mathcal{T}$, and $G=\gamma(F-Q)$ for some $\gamma\geq 0$. The converse is also true: If $F$ is optimal then $F=\Pi(Q)$ for some $Q$ with $-FA^{T}\in\mathcal{T}$ and $G=\gamma(F-Q)$ for some $\gamma\geq 0$. (Indeed, this holds with $\gamma=1$ and $Q=F-G$.) This optimality condition is readily interpreted: It states that $F$ is a fixed point of a projected gradient step with step size $\gamma$.
@@ 1.3 Related work
Specialized to GPU-based optimizers for network flow optimization, [WZR+18] considers implementing a parallel routing algorithm on GPUs for SDN networks, which solves the Lagrangian relaxation of a mixed integer linear program. [KOY+15] implements a genetic method on GPUs for solving an integer linear program formulation of routing problem. [ZAC23] considers a linear program formulation of multicommodity network flow problems and constructs a deep learning model for generating new columns in delayed column generation method. [WHH12] implements an asynchronous push-relabel algorithm for single commodity maximum network flow problem, which is CPU-GPU hybrid. [LHQ+24] exploits exactly the same flow aggregation formulation of MCF following [YDLB19] as we do and trains a neural network model for minimizing unconstrained Lagrangian relaxation objective, and feeds the result as warm start to Gurobi [Gur24] to get the final answer. [YP24] integrates a source-based flow aggregation formulation of the multicommodity flow problem into solving the combined transportation model and exploits an accelerated variant of proximal alternating predictor-corrector algorithm. The authors claim that the proposed algorithm is GPU-friendly, but the numerical experiments are CPU-based, and involve small size networks. [KYP+23] adopts a primal-dual gradient method for solving combined traffic models, which however is not GPU-oriented.
@@ 2.6 Implementation details
Some recent work develops a more efficient method to compute the projection onto the simplex set ([DSSSC08] and [Con16] for example); we adopt the simpler algorithm described above for implementation simplicity.
@@ Acknowledgements
We thank Anthony Degleris and Parth Nobel for valuable discussions on implementation details. We also thank Demyan Yarmoshik for very useful feedback for the revision of our original manuscript.